# Insight-HXMT View of the BHC Swift J1727.8-1613 during its outburst in 2023

### v2

@@ Abstract
The transient Galactic black hole candidate Swift J1727.8-1613 went through an outburst for the very first time that started in August 2023 and lasted for almost 6 months. We study the timing and spectral properties of this source using publicly available archival Insight-HXMT data for the first 10 observation IDs that last from MJD 60181 to 60198 with a total of 92 exposures for all three energy bands. We extracted the quasi-periodic oscillation properties by model fitting the power density spectrum and from those properties we designate that the QPOs are type-C in nature. We also conclude that the origin of the QPOs could be the shock instabilities in the transonic advective accretion flows around black holes. The spectral analysis was performed using simultaneous data from the three on-board instruments LE, ME, and HE of Insight-HXMT in the broad energy band of $2-150$ keV. To achieve the best fit, spectral fitting required a combination of models e.g. interstellar absorption, power-law, multi-color disk-blackbody continuum, gaussian emission/absorption, and reflection by neutral material. From the spectral properties, we found that the source was in an intermediate state at the start of the analysis period and was making a transition toward the softer states. The inner edge of the accretion disk moved inward in progressive days following the spectral nature. We found that the source has a high inclination. The average hydrogen column density estimated from the model fitting is $0.27_{-0.17}^{+0.08}\times 10^{22}$ cm-2.
X-rays: binaries – stars: black holes – stars: individual (Swift J1727.8-1613) – accretion, accretion disk – shock – radiation
journal: AAS
@@ 1 Introduction
Stellar-mass black holes (SBHs) are one of the end products of the death of massive stars. They reside in binary systems where a companion star supplies matter and during accretion they emit radiation. A significant amount of their radiation comes out in the form of X-rays, therefore, they are called black hole X-ray binaries (BHXRBs). Depending on the mass of the companion star, they can be mainly of two types: low-mass X-ray binaries (LMXRBs) or high-mass X-ray binaries (HMXRBs). The LMXRBs consist of type-A or later type of stars, whereas the HMXRBs consist of giant O or B-type of stars (White et al. 1995; Tetarenko et al. 2016). During the transfer of mass from the companion, matter accumulates at a distance from the BH due to insufficient viscosity to drive the matter inward, which is called the pile-up radius (Chakrabarti et al. 2019; Chatterjee et al. 2022). Due to the accumulation of more matter, the temperature will rise. The gradual increase in temperature develops instability at the pile-up radius, which further increases the viscosity. At some point in time when the accumulated matter gains enough viscosity to push forward, accretion starts. On the other hand, irradiation of the accumulated matter at the outer boundary by the central object can also trigger mass accretion when the temperature of the accumulated matter crosses a critical limit (King & Ritter 1998; Mondal 2020). Thus, these systems make ideal laboratories for studying the physics of accretion. The BHXRBs mostly stay in a quiet dormant state when their flux remains at a very low detection level. Episodically, the radiation level increases by a very large factor and they become easily detectable, which is known as an outburst (Remillard & McClintock 2006). The X-ray luminosity increases by several orders of magnitude during an outburst when compared to a quiescence. Depending on the nature of the outburst, the sources are classified into two types: persistent and transient sources. Most of the BHXRBs are transient type of sources. They stay in the dormant state for most of their lifetimes. The luminosity becomes $\sim 10^{35}$ erg s-1 during episodic outbursting phase.
@@
During the onset of an outburst, the flux in the light curve changes noticeably. Based on its nature, the outbursts are divided into two types (Debnath et al. 2010): fast rise slow decay (FRSD) and slow rise slow decay (SRSD). However, there are also some other classified outburst types, e.g., according to Zhang et al. (2019), outbursts can also be classified into glitch, reflare, multipeak, and mini-outburst. There are generally four designated spectral states during a complete BH outburst (Remillard & McClintock 2006). They are known as the hard state (HS), hard intermediate state (HIMS), soft intermediate state (SIMS), and soft state (SS). Generally, when an outburst starts, it starts in the HS. Then it slowly moves toward the SS through the HIMS and SIMS (Belloni et al. 2005, 2011, and references within). After it reaches its SS, the rising phase generally comes to an end and the outburst starts its decaying phase. In the decaying phase, it goes in the opposite direction and reaches the declining HS. When a BH outburst goes through all these states, it is known as a type-I outburst. When the soft state is absent, it is known as a ‘failed’ or type-II outburst.
@@
The radiation spectrum of a BHXRB generally consists of two main components. One is the soft thermal multi-colour blackbody component, which is due to the blackbody radiation of the seed photons in the standard accretion disk (Shakura & Sunyaev 1973; Novikov & Thorne 1973). The other is the hard power-law component, which can extend up to very high energies. The origin of this component is due to the inverse Comptonization of a fraction of seed photons from the disk, which is intercepted by a hot Compton cloud (Sunyaev & Titarchuk 1980, 1985; Haardt & Maraschi 1993; Zdziarski et al. 1993; Titarchuk 1994; Chakrabarti & Titarchuk 1995; Życki et al. 1999). When the outburst is in HS, there is a very small contribution from the thermal blackbody component and mostly the (non)thermal powerlaw component dominates. In SS, the thermal blackbody component becomes dominant while in the intermediate states (HIMS and SIMS), the contribution from the two components stays comparable. Other than these two components, in the case of a high soft state (HSS), there could be the presence of bulk motion Comptonization (BMC) which takes place due to the relativistic speed of matter when it reaches a closer to the BH (Blandford & Payne 1981a,b; Payne & Blandford 1981; Chakrabarti & Titarchuk 1995; Psaltis & Lamb 1997; Borozdin et al. 1999; Psaltis 2001). It was recently observed by Chatterjee et al. (2023) that BMC was present in the radiation spectrum of the black hole candidate (BHC) MAXI J0637-430.
@@
In addition, reprocessing of the Comptonized photons from the Compton cloud by the accretion disk is known as reprocessed radiation (George & Fabian 1991; Ross & Fabian 2005; Garcia & Kallman 2010), which generates reflection spectra. As a result, the Fe $K\alpha$ emission line at $\sim 6.4$ keV and a reflection hump above 20 keV may originate. Therefore, the presence of complex features in the spectrum in BHXRBs makes them important to study to understand their origin. This can be achieved by fitting broadband data using multicomponent models. In this regard, high-resolution broadband data of Insight-HXMT (Zhang et al. 2020) can be very useful.
@@
Besides rich spectral features and variabilities, temporal properties are equally important to understanding the dynamics of the accreting gas around the BHs. During an outburst, it has been observed that the light curves show very small timescale variabilities, especially in the high-energy bands. The Fourier transformation of the light curve imprints such variabilities as both broadband noise and narrow features in the power density spectrum or PDS (van der Klis 1989). The broadband noise is spread over a large frequency range modeled by power-law function. The narrow feature is a power peak in narrow frequency ranges, known as the quasi-periodic oscillation (QPO), modeled by one or multiple Lorentzian profiles. In BHXRBs, the low-frequency QPOs (LFQPOs) are often observed mainly due to their geometrical origin. Depending on their properties like the frequency ($\nu$), $Q$-value ($=\nu/\delta\nu$, where $\delta\nu$ is the full width at half maximum or FWHM), (%)RMS, etc., LFQPOs are classified into three different types: type A, B, and C (Casella et al. 2005). Although their origin is still a topic of debate, several ideas were put forward e.g., Lense-Thirring precession (Stella et al. 1999; Ingram et al. 2009), magneto-acoustic waves (Titarchuk et al. 1998), accretion-ejection instability (Tagger & Pellat 1999), shock oscillation model (Molteni et al. 1996; Chakrabarti et al. 2005, 2008, 2015). The shock oscillation model can explain both the spectral and temporal properties simultaneously, which makes it a more comprehensive model.
@@
As both the properties are coming from the same system, and changes in spectral states can also dictate the types of QPOs, it is believed that both are interlinked. A good correlation is observed between the two using just the light curve properties, e.g., the hardness ratio or HR, hardness intensity diagram, or HID (Homan et al. 2001). Some model-fitted approaches also showed that the changes in both properties are due to the interplay between different mass accretion rates or variation in viscosity, can be understood through accretion rate ratio intensity diagram or ARRID (Mondal et al. 2014; Jana et al. 2016; Chatterjee et al. 2020). The RMS-intensity diagram or RID (Munoz-Darias et al. 2011), and hardness ratio-intensity diagram or HRD (Belloni et al. 2005) can also make a link between spectral and temporal properties from pure observational ground.
@@
The BHC Swift J1727.8-1613 was discovered very recently on 24 August 2023 or MJD 60180 (Kennea & Swift Team 2023). MAXI (Matsuoka et al. 2009) detected the source to be very bright with a flux of $7$ Crab in the $2-20$ keV energy range. Since then multi-wavelength follow-up observations were carried out by several ground-based and space-bourne telescope facilities (Baglio et al. 2023; Miller-Jones et al. 2023; Negoro et al. 2023b; O’Connor et al. 2023; Wang & Bellm 2023; Williams-Baldwin et al. 2023). The source is reported to be located at a distance of $2.7\pm 0.3$ kpc (MataSanchez et al. 2024) and a mass of $10\pm 2~{}M_{\odot}$ (Svoboda et al. 2024). The revelation of the presence of hydrogen and helium emission lines in the optical spectrum led to its classification as an LMXRB (Castro-Tirado et al. 2023). The X-ray spectrum and other properties point to its nature as a BHXRB (Liu et al. 2023; Sunyaev et al. 2023). This was supported strongly after the detection of type C QPOs (Palmer & Parsotan 2023; Draghis et al. 2023; Bollemeijer et al. 2023; Mereminskiy et al. 2023) and a flat radio spectrum (Miller-Jones et al. 2023; Bright et al. 2023). There is X-ray monitoring of the source by NICER, NuSTAR, Insight-HXMT, IXPE, etc. Using IXPE data on September 7, 2023, Veledina et al. (2023) estimated its polarization with a polarization degree (PD) of $4.1\%\pm 0.2\%$ and polarization angle (PA) of $32^{\circ}.2\pm 1^{\circ}.$3. There was a further polarization study of the source by Ingram et al. (2023).
@@
At the starting of the outburst, the spectral nature was consistent with a hard state with a power-law photon index ($\Gamma$) $\sim 1.5-1.7$ (Liu et al. 2023). Later, using the NICER observations, Bollemeijer et al. (2023) reported the softening of the spectrum with a substantial contribution of the soft blackbody component after 25 days from the starting of the outburst. The presence of LFQPOs was reported by Draghis et al. (2023) and the evolution of LFQPOs at the high energy band using AstroSat data was reported by Nandi et al. (2024). The source also showed soft time lag of $0.014\pm 0.001$ sec between energy bands of $3-10$ and $0.5-3$ keV on August 29 2023 (MJD 60185) (Debnath et al. (2023) using NICER data). With reflection spectroscopy, authors found the disk inclination to be $\sim 87^{\circ}$. However, to date, there is no detailed study on the origin of the spectral and temporal properties of the source, which prompted us to analyze the broadband X-ray data from Insight-HXMT.

### v3

@@ Abstract
The transient Galactic black hole candidate Swift J1727.8-1613 went through an outburst for the very first time in August 2023 and lasted for almost 6 months. We study the timing and spectral properties of this source using publicly available archival Insight-HXMT data for the first 10 observation IDs that last from MJD 60181 to 60198 with a total of 92 exposures for all three energy bands. We have detected quasi-periodic oscillations (QPOs) in a frequency range of $0.21\pm 0.01$ $1.86\pm 0.01$ Hz by fitting the power density spectrum. Based on the model-fitted parameters and properties of the QPOs, we classify them as type-C in nature. We also conclude that the origin of the QPOs could be the shock instabilities in the transonic advective accretion flows around black holes. The spectral analysis was performed using simultaneous data from the three on-board instruments LE, ME, and HE of Insight-HXMT in the broad energy band of $2-150$ keV. To achieve the best fit, spectral fitting required a combination of models e.g. interstellar absorption, power-law, multi-color disk-blackbody continuum, Gaussian emission/absorption, and reflection by neutral material. From the spectral properties, we found that the source was in an intermediate state at the start of the analysis period and was transitioning to the softer states. The inner edge of the accretion disk moved inward in progressive days following the spectral nature. We found that the source has a high inclination of $78^{\circ}-86^{\circ}$. The hydrogen column density from the model fitting varied in the range of $(0.12\pm 0.02-0.39\pm 0.08)\times 10^{22}$ cm-2.
X-rays: binary stars (1811); black holes (162); Stellar accretion disks (1579); Shocks (2086); Compact radiation sources (289)
@@ 1 Introduction
Stellar-mass black holes (SBHs) are one of the end products of the death of massive stars. They reside in binary systems where a companion star supplies matter and during accretion they emit radiation. A significant amount of their radiation comes out in the form of X-rays, therefore, they are called black hole X-ray binaries (BHXRBs). Depending on the mass of the companion star, they can be mainly of two types: low-mass X-ray binaries (LMXRBs) or high-mass X-ray binaries (HMXRBs). The LMXRBs consist of type-A or later type of stars, whereas the HMXRBs consist of giant O or B-type of stars (e.g., White et al. 1995; Tetarenko et al. 2016). During the transfer of mass from the companion, matter accumulates at a distance from the BH due to insufficient viscosity to drive the matter inward (Mineshige 1993; Kuulkers et al. 1997; Parmar et al. 1997; Lasota 2001), which is called the pile-up radius (Chakrabarti et al. 2019; Chatterjee et al. 2022). Due to the accumulation of more matter, the temperature of the disk rises. The gradual increase in temperature develops instability at the pile-up radius, which further increases the viscosity. At some point in time when the accumulated matter gains enough viscosity to push forward, accretion starts. On the other hand, irradiation of the accumulated matter at the outer boundary by the central object can also trigger mass accretion when the temperature of the accumulated matter crosses a critical limit (King & Ritter 1998; Mondal 2020). Thus, these systems make ideal laboratories for studying the physics of accretion. BHXRBs mostly stay in a quiet dormant state when their flux remains at a very low detection level. Episodically, the radiation level increases by a very large factor and they become easily detectable, which is known as an outburst (Tanaka & Shibazaki 1996; Remillard & McClintock 2006). The X-ray luminosity increases by several orders of magnitude during an outburst when compared to a quiescence. Depending on the nature of the outburst, the sources are classified into two types: persistent and transient sources. Most of the BHXRBs are transient type of sources. They stay in the dormant state for most of their lifetimes. The luminosity becomes $\sim 10^{37-38}$ erg s-1 during episodic outbursting phase (Tanaka & Shibazaki 1996).
@@
During the onset of an outburst, the flux in the light curve changes noticeably. Based on its nature, the outbursts are divided into two types (Debnath et al. 2010): fast rise slow decay (FRSD) and slow rise slow decay (SRSD). However, there are also some other classified outburst types, e.g., according to Zhang et al. (2019), outbursts can also be classified into glitch, reflare, multipeak, and mini-outburst. There are generally four designated spectral states during a complete BH outburst (Remillard & McClintock 2006). They are known as the hard state (HS), hard intermediate state (HIMS), soft intermediate state (SIMS), and soft state (SS). Generally, when an outburst starts, it starts in the HS. Then it slowly moves toward the SS through the HIMS and SIMS (Belloni et al. 2005, 2011, and references within). After it reaches its SS, the rising phase generally comes to an end and the outburst starts its decaying phase. In the decaying phase, it goes in the opposite direction and reaches the declining HS. When a BH outburst goes through all these states, it is known as a type-I outburst. When the soft state is absent, it is known as a ‘failed’ or type-II outburst (Hynes et al. 2000; Brocksopp et al. 2001; Belloni et al. 2002; Brocksopp, Bandyopadhyay & Fender 2004; Capitanio et al. 2009; Curran & Chaty 2013; Del Santo et al. 2016; Tetarenko et al. 2016; Garcia et al. 2019; Alabarta et al. 2021).
@@
The energy spectrum of a BHXRB generally consists of two main components. One is the soft thermal multi-colour blackbody component, which is due to the blackbody radiation of the seed photons in the standard accretion disk (Shakura & Sunyaev 1973; Novikov & Thorne 1973). The other is the hard power-law component, which can extend up to very high energies. The origin of this component is thought to be due to the inverse Comptonization of a fraction of seed photons from the disk, which is intercepted by a hot Compton cloud (Sunyaev & Titarchuk 1980, 1985; Haardt & Maraschi 1993; Zdziarski et al. 1993; Titarchuk 1994; Chakrabarti & Titarchuk 1995; Życki et al. 1999). When the outburst is in the HS, there is a very small contribution from the thermal blackbody component and mostly the (non)thermal powerlaw component dominates. In the SS, the thermal blackbody component becomes dominant while in the intermediate states (HIMS and SIMS), the contribution from the two components stays comparable. Other than these two components, in the case of a high soft state (HSS), there could be the presence of bulk motion Comptonization (BMC) which takes place due to the relativistic speed of matter when it reaches a closer to the BH (Blandford & Payne 1981a,b; Payne & Blandford 1981; Chakrabarti & Titarchuk 1995; Psaltis & Lamb 1997; Borozdin et al. 1999; Psaltis 2001). It was recently observed by Chatterjee et al. (2023) that BMC was present in the radiation spectrum of the black hole candidate (BHC) MAXI J0637-430.
@@
In addition, reprocessing of the Comptonized photons from the Compton cloud by the accretion disk is known as reprocessed radiation (George & Fabian 1991; Ross & Fabian 2005; Garcia & Kallman 2010), which generates reflection spectra. As a result, the Fe $K\alpha$ emission line at $\sim 6.4$ keV and a reflection hump above 20 keV may originate. The presence of these complex features in the spectrum of BHXRBs makes it interesting to study. It is important to study and model these features to understand their origin. This can be achieved by fitting broadband data using multicomponent models. In this regard, high-resolution broadband data of Insight-HXMT (Zhang et al. 2020) can be very useful.
@@
Besides rich spectral features, temporal properties are equally important to understanding the dynamics of the accreting gas around the BHs. During an outburst, it has been observed that the light curves show changes in very small timescale, especially in the high-energy bands. The Fourier transformation of the light curve imprints such changes as both broadband noise and narrow features in the power density spectrum or PDS (van der Klis 1989). The broadband noise is spread over a large frequency range modeled by power-law or broad Lorentzian functions. The peak-like feature is a power peak in narrow frequency ranges, known as the quasi-periodic oscillation (QPO). The PDS is modeled by one or multiple Lorentzian profiles along with the power-law model. The Lorentzian model helps estimate the properties of the QPO (Nowak 2000; Belloni et al. 2002). Low frequency quasi periodic oscillations (LFQPOs) lie in the frequency range of a few mHz to $\sim 30$ Hz (Belloni et al. 2002; Casella et al. 2005; Motta et al. 2015). Depending on their properties, e.g., the frequency ($\nu$), $Q$-value ($=\nu/\delta\nu$, where $\delta\nu$ is the full width at half maximum or FWHM), (%)RMS, etc., LFQPOs are classified into three different types: type A, B, and C (Casella et al. 2005). Several ideas were put forward to explain the LFQPO phenomenon in BHs. Although type-C LFQPO is now a well-studied phenomenon, there is still debate about its origin, as several models can explain it. Some of the often used explanations are the Lense-Thirring precession (Stella et al. 1999; Ingram et al. 2009), magneto-acoustic waves (Titarchuk et al. 1998), accretion-ejection instability (Tagger & Pellat 1999) and the shock oscillation model (Molteni et al. 1996; Chakrabarti et al. 2005, 2008, 2015). However, compared to type-C, the other two types’ origin is still not clearly known.
The shock oscillation is one of the features of the two-component advective flow (TCAF; Chakrabarti & Titarchuk 1995) model. This model suggests that there are two different forms of angular momentum distribution in matter supplied from the companion star, Keplerian and sub-Keplerian. Due to its high viscosity, the Keplerian component moves in a viscous timescale and forms a thin accretion disk on the equatorial plane. The accretion disk moves inwards as this matter gains critical viscosity. Less viscous accretion flow characterizes the sub-Keplerian one, which falls radially in a free fall timescale. This matter resides above and below the Keplerian disk. The sub-Keplerian component, which is optically thin, creates a shock front where infalling matter virtually stops because of the tug of war between centrifugal and gravitational forces. The boundary layer, sometimes referred to as the CENBOL (centrifugal pressure-supported boundary layer, Chakrabarti & Titarchuk 1995), is the shock front. In this model, the post-shock region is the ‘Compton’ cloud region, the repository of hot electrons. As blackbody photons are produced by the disk, the soft multicolor blackbody component can be explained by the Keplerian component. Some of the captured soft disk photons are upscattered and released as hard power-law photons by the CENBOL, which serves as the Comptonizing region. During the start of an outburst, the CENBOL forms farther away from the BH, and the disk is also truncated at a very large radius. As a result, photons originating from the disk become hard through inverse Compton scattering by hot electrons, and thus we always observe a hard state during the onset of an outburst from BHs. As time progresses, the Keplerian disk moves inwards, producing more blackbody radiation. More soft photons are intercepted in the CENBOL, which are upscattered and help to cool down the CENBOL. More photons are emitted in the whole process and flux increases and the source goes toward a softer state. In a high soft state, the CENBOL is totally quenched, and the disk reaches very close to the black hole. Thus, only the disk photons now contribute to the radiation, which makes the spectrum soft.
Besides describing the spectral features and their changes, this model also explains the QPO properties. In the HS, when the CENBOL is bigger, it could produce LFQPOs due to the resonance oscillation of the shock, when the infall and cooling timescales match (Molteni et al. 1996, Chakrabarti et al. 2015, Garain et al. 2014). As the outburst progresses, CENBOL loses its size, and we see a rise in the frequency of the QPO. This is discussed with mathematical relation in the discussion section. As the SS approaches, the CENBOL shrinks in size, and there is no oscillation, cooling takes over the heating timescale. Thus, we do not see any QPO in the SS. In between the HS and SS, when the CENBOL is intermediate in size, there could be the presence of either type-C or type-A/B QPOs. In the HIMS, we see type-C, whereas, in the SIMS, mostly type-A/B is seen.
@@
As changes in both the spectral and timing properties come from the same system and changes in spectral states can also dictate the types of QPOs, there could be a possible correlation between the spectral and temporal properties. A connection between these two properties is observed using a hardness intensity diagram, or HID (Homan et al. 2001), which shows the variation of the flux in the light curve (considered here as timing property) with the hardness ratio, or HR, which shows roughly the variation of the spectral nature of the outburst (considered as spectral property here). A detailed outburst study showed that the changes in both properties are due to the interplay between different mass accretion rates or variation in viscosity, which can be understood through the accretion rate ratio intensity diagram or ARRID (Mondal et al. 2014; Jana et al. 2016; Chatterjee et al. 2020). The RMS-intensity diagram (RID, Munoz-Darias et al. 2011), and hardness ratio-intensity diagram or HRD (Belloni et al. 2005) can also link spectral and temporal properties from a pure observational ground. As shown in Chakrabarti et al. (2015), when the cooling timescale is within 50% of the infall/heating timescale, the QPO originates, which is in accord with the spectral model fitted ARRID in Mondal et al. (2014), where QPOs disappeared for ARR $<0.5$ in the ARRID. Unlike most of the explanations in the literature, the shock oscillation model in TCAF paradigm is successful in addressing the spectral and temporal properties responsible for the origin of QPOs, which prompted us to consider this model for rest of the paper to explain QPOs observed in Swift J1727.8-1613 during Insight-HXMT era.
@@
The BHC Swift J1727.8-1613 was discovered very recently on 24 August 2023 or MJD 60180 (Kennea & Swift Team 2023). MAXI (Matsuoka et al. 2009) detected the source to be very bright with a flux of $7$ Crab in the $2-20$ keV energy range. Since then multi-wavelength follow-up observations were carried out by several ground-based and space-bourne telescope facilities (Baglio et al. 2023; Miller-Jones et al. 2023; Negoro et al. 2023b; O’Connor et al. 2023; Wang & Bellm 2023; Williams-Baldwin et al. 2023). The source is reported to be located at a distance of $2.7\pm 0.3$ kpc (MataSanchez et al. 2024) and a mass of $10\pm 2~{}M_{\odot}$ (Svoboda et al. 2024). The revelation of the presence of hydrogen and helium emission lines in the optical spectrum led to its classification as an LMXRB (Castro-Tirado et al. 2023). The X-ray spectrum and other properties point to its nature as a BHXRB (Liu et al. 2023; Sunyaev et al. 2023). This was supported strongly after the detection of type C QPOs (Palmer & Parsotan 2023; Draghis et al. 2023; Bollemeijer et al. 2023; Mereminskiy et al. 2023) and a flat radio spectrum (Miller-Jones et al. 2023; Bright et al. 2023). There have been X-ray monitoring of the source with NICER (O’Connor et al. 2023), NuSTAR (Dovciak et al. 2023), Insight-HXMT (Peng et al. 2024), IXPE (Veledina et al. 2023), AstroSat (Katoch et al. 2023). Using IXPE data on September 7, 2023, Veledina et al. (2023) estimated its polarization with a polarization degree (PD) of $4.1\%\pm 0.2\%$ and polarization angle (PA) of $32^{\circ}.2\pm 1^{\circ}.$3. There was a further polarization study of the source by Ingram et al. (2023).
@@
At the begining of the outburst, the source was consistent with being in the hard state, described by a power-law with a photon index ($\Gamma$) $\sim 1.5-1.7$ (Liu et al. 2023). Later, using the NICER observations, Bollemeijer et al. (2023) reported the softening of the spectrum with a substantial contribution of the soft blackbody component after 25 days from the start of the outburst. The presence of LFQPOs was reported by Draghis et al. (2023) and the evolution of LFQPOs at the high energy band using AstroSat data was reported by Nandi et al. (2024). The source also showed soft time lag of $0.014\pm 0.001$ sec between energy bands of $3-10$ and $0.5-3$ keV on August 29 2023 (MJD 60185) (Debnath et al. (2023) using NICER data). With reflection spectroscopy, these authors found the disk inclination to be $\sim 87^{\circ}$. However, none of the studies in the literature discussed the origin of the spectral and temporal properties in the context of the shock oscillation model, which prompted us to analyze the broadband X-ray data from Insight-HXMT and interpret them in light of accretion with shocks.